Solve the following equation:
[-3 -1 -71] = [-3 41 7] - 3x

Answer by jsmallt9(3758)

[-3 -1 -71] = [-3 41 7] - 3x
Subtract [-3 41 7] from each side:
[0 -42 -78] = -3x
Multiply each side by -1/3:
(-1/3) = x
[0 14 26] = x